The weekly salaries of teachers in one state are normally distributed with a mean of 560.0 dollars and a standard deviation of 39.0 dollars. What is the probability that a randomly selected teacher earns more than 605.0 a week?The weekly salaries of teachers in one state are normally distributed with a mean of 560.0 dollars and a standard deviation of 39.0 dollars. What is the probability that a randomly selected teacher earns more than 605.0 a week? SELECT ALL APPLICABLE CHOICES А) B) 87.57187% 14.42813% D) 8.428135% 4.428135% E) F) None of These 12.42813%

My answer, B, is not correct.
Answer #1

\mu = 560

\sigma = 39

To find P(X>605):
Z = (605 - 560)/39 = 1.1538

By Technology, Cumulative Area Under Standard Normal Curve = 0.8757187

So,

P(X>605) = 1 - 0.8757187 = 0.1242813 = 12.42813 %

So,

Correct option:

E) 12.42813